Operations Bulletin 160315

From GridPP Wiki
Jump to: navigation, search

Bulletin archive

Week commencing 9th March 2015
Task Areas
General updates

Tuesday 10th March

  • CERN computing seminars.
  • Status of WNs in CVMFS
  • Suggestion to write actions to Vidyo chat window
  • Problem with dashboards – following announced GOCDB changes last week.
  • cvmfs 2.1.20 is currently being certified at CERN and RAL. A small update to cvmfs-puppet is required to deploy cvmfs 2.1.20 clients.
  • A new version of the CA Portal was released last week. It supports email address updates.
  • JK requests those with host certificates with embedded email addresses to phase them out.
  • CERN accounts: Note WLCG office acting as CERN guarantor for everyone, so by default everyone will be”‘LCG"’d in the database (new HR rules mean Externals automatically take on the organic-unit of the guarantor).
  • RAL T1: Dell Force10 Z9000 Firmware update. Was on version running on BIOS and boot code Testing version is which runs on the same BIOS but requires boot code
  • For LIGO, Rob Quick informs JC that OSG has recently re-engaged with the VO and will update us on their VOMS situation in the next week.
  • ARC CONDOR CPU time fix - see Steve Jones's TB-S email on 3rd March. (related to ATLAS submitting some longer running Multi-Core jobs which get killed with the original configuration?)
  • A reminder that on Tuesday 10th there is a pre-GDB: agenda. AM: HEP and other sciences. PM: Cloud issues.
  • Wednesday 11th there is a GDB: Agenda covering EGI/WLCG work, European procurements, WLCG operational costs and supporting other sciences.

Tuesday 3rd March

  • There was an EGI OMB last Thursday. Highlights for awareness:
    • EGI core activities to be reviewed for next phase.
    • EGI provides a catch-all CA; Request to register iRODS endpoints in GOCDB; EU Brazil Cloud Connect mentioned; May conference registrationopen. Next OMB 26th March.
    • Check the status of multi-core publishing.
    • New release of ops portal coming 16th March. See Schedule and version 3.1.2 changes.
    • New performance documents. QoS; ROD; Availability/Reliability.
    • Security: FedCloud survey done. EGI-CSIRT collaboarating with FedCloud. Security challenges coming - communications one done; next pull in payload and assess capabilities. Also looking at VM incident handling.
    • ARGO monitoring: Modular architecture. Example deployment models given for EGI/VO/NGI. Uses message brokering - between collectors and review engine. Listed advantages over ACE. Problems with SAM Nagios given (e.g. effort to upgrade, support). ARGO is SAM refactored (no heavy DB). -> Need a UK view on this as being pushed for EGI-Engage.
    • VO SLAs: to clarify expectations. Some updates including FedCloud inclusion.
    • EGI Engagement (strategy): Call for unfunded partners to join core competency centres: BBMRI, DARIAH, EISCAT_3D, ELIXIR, EPOS, INSTRUCT, LifeWatch & Disaster mitigation.
    • New infrastructures in preparatory project phase: MIRRI – microbal resources; EuroDISH – food, nutrition, diseases; ISBE – systems biology. Opportunities for national engagement: EMBRC: Marine biological resource centre (... + UK); ERINHA: Highly Pathogenic Agents (...+ UK); Euro-BioImaging (...+ UK & EMBL).INSTRUCT: Structural biology (many countries). CLARIN (call out).
    • Genome analysis & protein folding pilot ongoing: Chipster, RSAT, READemption.
    • Ongoing integration of ELIXIR reference datasets into EGI. Several use-cases for FedCloud mentioned (any more sites want to join?)
  • Preliminary agendas are available for the pre-GDB and GDB.
  • The deadline for CHEP registrations is 15th March. (Bulletins)

Tuesday 24th February

WLCG Operations Coordination - Agendas

Monday 9th March

  • There was a WLCG ops meeting on Thursday: Agenda | Minutes.
  • In summary:
    • News: Looking at survey – GDB & Workshop.
    • M/W news: Baselines – new FTS 3.3.32 (shares fix).
    • M/W issues: RAL problem upgrading gfal2 in WN ( conflict with gfal). New vulnerability (SSL).
    • T0/T1 services: Various FTS and SE upgrades at CERN, RAL and BNL...
    • Question about WN and UI status and maintenance wrt CVMFS.
    • T0 news: LHCb LFC migration fine. Checking who uses LFC. VOMS-admin in place - some issues require follow-up.
    • T1/T2: NTR
    • ALICE: High activity. T1/T2 workshop. Some VOMS-admin problems.
    • ATLAS: Taking cosmics. MC15 started (consider start of Run2). Care with maintenance requested. Transatlantic link problems encountered.
    • CMS: Taking cosmics. Moderate load. Tier-1 tape tests continue. VOMRS migration in progress. Migration to a single global Condor pool for Analysis and Production (For T2s: 80% VOMS pilot; 10% VOMS production; 10% other).
    • LHCb: "Run1 Legacy stripping" finished. Some VOMRS migration issues. LFC to DFC migration done.
    • glexec: Panda campaign at 63 sites (+9)
    • RFC proxies: Presentation at GDB. Checking experiment usage. SAM-Nagios fix easy. Default later this year.
    • MJ features: NTR
    • M/W readiness: Recent Storm/dCache/DPM validations done, more happening. ARC-CE tests for CMS in preparation. Next meeting 18th March.
    • Multicore: Main objectives achieved. Entering 'passive' mode.
    • IPv6: NTR
    • Squid proxy & http proxy discovery: NTR
    • Network & transfers WG: [(https://indico.cern.ch/event/372546/ Meeting 18th Feb]. Deadline for 3.4.1 passed. Campaign to correct some configurations. Testbed for 3.4.2rc. New meshes (inc. IPv6). Testing and evaluation of the pilot instances for esmond/maddash ongoing. Production instance: psomd.grid.iu.edu. LHCb pilot project to provide experiment agnostic prototype to access central datastore (esmond). Extending ATLAS FTS performance study to CMS and LHCb. Next meeting 18th March.
    • Temporary ATLAS solution has been found to publish the HTCondor CEs in the BDII and OIM in a way that satisfies both ATLAS and SAM needs.
    • Next meeting 19th March.

Monday 2nd March

  • There is a WLCG ops meeting coming up this Thursday. Any T1/T2 issues?
  • There was a perfSONAR data extraction meeting today.

Tier-1 - Status Page

Tuesday 10th March

  • We had a problem on our network around 3pm last Friday (6th March) with very high packet lost for around 20 minutes.
  • The problems with our primary network router are still being followed up. We had a problem and were not able to try out the spare switch last Thursday. This has been re-scheduled for early on this Thursday morning (12th March). An 'At Risk' has been declared in the GOC DB.
  • Tomorrow (Wednesday 11th March) we plan to update the firmware in our network switches to fix a problem of periodic reboots.
Storage & Data Management - Agendas/Minutes

Tuesday 9th March

  • DPM with puppet issues (trying this documented approach.
  • Command for backing up DPM (see email to TB-SUPPORT on 3rd) - is it documented!?

Tuesday 3rd March

  • DiRAC FS backup to RAL (and UCL) very likely to go ahead. GridPP encourages support from co-located sysadmins to help setup transfers... but first there needs to be a discussion of how the data will be handled. More soon.

Monday 23rd February

  • From February 17th expect xrootd version 4 will be added to the EPEL repository. Update not transparent if xrootd supported.

Tuesday 17th February

  • Chris suggests linking the WebDav page from the storage wiki.

Accounting - UK Grid Metrics HEPSPEC06 Atlas Dashboard HS06

Tuesday 10th March

  • Possible delay at Birmingham. Worth checking.

Tuesday 24th February

  • UKI-SOUTHGRID-SUSX now highlighted as not publishing for 3 months.
Documentation - KeyDocs

See the worst KeyDocs list for documents needing review now and the names of the responsible people.

Tuesday 3rd March

  • Proposals for about 4 keydocs to be removed/downgraded. For discussion at next core-ops or focus meeting.

Tuesday 24th February

Tuesday 17th February

  • KeyDocs review continues. Deadline for removals requests 25th February.

Tuesday 10th February

  • Survey of KeyDocs has started. If you own a document please take note. Documents remaining after this exercise will need to be brought fully up-to-date.
  • A reminder that Steve created this Current Activities capture page. Only ARGUS at the moment!?
  • See also the documents from Tom linked in the 'Other VOs' section.

Tuesday 3 Feb

  • New section in Wiki called "Project Management Pages".
The idea is to cluster all Self-Edited Site Tracking Tables
in here. Sites should keep entries in Current Activities
up to date. Once a Self-Edited Site Tracking Tables has
served its purpose, PM to move it to  Historical Archive 
or otherwise dispose of the table.
Interoperation - EGI ops agendas

Monday 9th March

  • The agenda for February's EGI ops meeting is here. Minutes are here
    • APEL 1.4.0
      • Added Month and Year columns to primary key of CloudSummaries table in cloud schema.
    • DPM-Xrootd 3.5.2 is in EPEL stable - this is the first version of the component compatible with xrootd4
    • gLExec-wn - v. 1.2.3: lcmaps-plugins-c-pep 1.3.0-1 & mkgltempdir 0.0.5-1
      • "The lcmaps-plugins-c-pep-1.3.0-1 preferably needs the argus-pep-api-c-2.3.0. This version will be released into EMI & UMD repositories in a near future."
    • UMD 3.11.0 released on 16.02.2014, UMD 3.11.1 released on 4.03.2014
    • lcg-CA 1.62 noted with an intention to broadcast these as they occur as opposed to monthly.
    • EGI looking at the decommissioning of SL5, possibly by end of 2015, as a byproduct of adding CentOS 7 to UMD. NGIs to make a note if extended SL5 support is required.
    • Vincenzo Spinoso has joined EGI Ops team from NGI_IT. Vincenzo will chair EGI Ops.
    • Next meeting is April 20th.

Monitoring - Links MyWLCG

Monday 7th December

On-duty - Dashboard ROD rota

Monday 9th March

  • There was a problem with the upgrade of the dashboard when it lost the scope of the ROD role.
  • Generally quiet.
  • Low availability alarm for RALPP.
  • Gareth->Kashif

Monday 23rd February

  • Another quiet week.
  • Gordon -> Andrew

Monday 16th February

  • Uneventful week. Nothing to report to ops meeting.
Rollout Status WLCG Baseline

Tuesday 17th February

  • A message from Maria Dimou: A request to you and the sites you collaborate with, to please install the MW Package Reporter.

The current version, documented in the wiki , it is very easy to install and addresses all site and security requirements.

  • Quick check on: 1. ARC CE readiness testing. 2. Machine job features testing.

Tuesday 20th January

  • From Cristina's GDB talk last week note that EMI repositories will be frozen and the product team releases will become UMD-preview (repository content and webpages similar but now managed).


Security - Incident Procedure Policies Rota

Monday 9th March

  • Trust anchor release 1.62 (missing announcement!).
  • From DG: "1.62 is indeed out and was released by the IGTF on Feb 23rd. The EGI release by an unfortunate coincidence got linked to a VOMS update for UMD (v3.3.3), so was released only today."

Tuesday 24th February

  • Tracking approvals in GOCDB.
  • EGI SVG Advisory - dCache vulnerability for some access methods [SVG EGI-SVG-2015-8183]
  • New security officer starts this week!

Tuesday 17th February

  • The IGTF is about to release (23/2) an update to the trust anchor repository (1.62)

Services - PerfSonar dashboard | GridPP VOMS

- This includes notifying of (inter)national services that will have an outage in the coming weeks or will be impacted by work elsewhere. (Cross-check the Tier-1 update).

Tuesday 10th March

  • From the recent WLCG meeting, two slides (1 & 2) give the direction of the network monitoring and metrics progress: integration of perfSONAR event types into experiment monitoring and an architecture for data to get from RSV probes to client. Components described on slide 3.
  • The next LHCOPN and LHCONE joint meeting will take place on Monday 1st and Tuesday 2nd of June 2015 in Berkeley (US) (hosted by LBL and ESnet).

Tuesday 3rd March

Tuesday 3rd February

Tuesday 20th January


Monday 9th March 2015, 15.00 GMT
From last week's crusty ticket round up:

Tier 1
109694- 28/10/14
Matt M has managed to reclaim his tickets after a certificate change orphaned his old ones. Progress has resumed. Duncan has asked Matt to retry his failing tests with a simple copy to local disk example.

108944- 1/10/14
CMS AAA access at RAL. Andrew posed a question to the CMS xroot experts last week - if you have their details it might be a good idea to involve them in the ticket.

108356- 10/9/14
This VMCatcher ticket is still stuck waiting for a reply. Deafening silence for our 100IT colleagues.

97485 - 21/9/13
The Jet LHCB ticket is in the state of being wrapped up. LHCB have been removed from the local configs, and the site has been removed from LHCB's. I believe that this ticket can be terminated.

110353 - 25/11/14
Dan's managed to get webdav working on one SE, but not t'other. Very strange, but Dan is investigating (see also 111942).

No movement on the three glexec tickets (none expected on the two tarball ones in the last week though), the Lancaster perfsonar ticket is still waiting on another batch of local tweaks (and I still need to make sense of what I'm seeing). Matt RB closed Sussex's perfsonar ticket though - nice one.

The "Normal" tickets:

Atlas gLexec Hammercloud failures (RALPP and Tier 1)
111699 (Tier 1)
111703 (RALPP)
These tests were waiting on a new stable job release being made - this has been delayed (hopefully out tomorrow).

This LHCB ticket about stalled jobs looks like it can be closed (LHCB no longer see a problem). Update - set to solved, the jobs were being killed for using too much memory.

A CMS user saw glexec failures on some nodes - Kashif asked for some more information but there has been no reply. I'd consider giving the user till the end of the week then closing the ticket if there's still no word. Waiting for reply (27/2)

Tools - MyEGI Nagios

Tuesday 17th February

  • Another period where message brokers were temporarily unavailable seen yesterday. Any news on the last follow-up?

Tuesday 27th January

  • Unscheduled outage of the EGI message broker (GRNET) caused a short-lived disruption to GridPP site monitoring (jobs failed) last Thursday 22nd January. Suspect BDII caching meant no immediate failover to stomp://mq.cro-ngi.hr:6163/ from stomp://mq.afroditi.hellasgrid.gr:6163/

VOs - GridPP VOMS VO IDs Approved VO table

Monday 9th March

  • SIXT VOMS parameters have changed.

Tuesday 17th February Two changes to approved VOs (https://www.gridpp.ac.uk/wiki/GridPP_approved_VOs)

  • LSST uses port 15003 (had been 15002, clashing with dzero)
  • t2k has included a note that it';s software is now distrinuted via CVMFS.

Tuesday 17th February

  • Some interest from UK LIGO users. Catalin setting up CVMFS. Issue with VOMS. "Our users have access to CILogon basic CA certificates, which I understand can in principal be entered into VOMS to provide a VO identity."

Tuesday 10th February

  • There is now a Quick Guide to CVMFS.
  • Part of Tom's quick guide series! There is also a Quick Guide to DIRAC.
  • Feedback on both welcome... on the CVMFS front some mapping issues for regional VOs are being examined.
  • GalDyn (UCLan galaxy simulations): test jobs have been submitted via DIRAC; now moving onto data storage testing via the official DIRAC tutorials.
  • CERN@school has its first student on the grid with a grid certificate! LUCID data arriving.
  • The CERN VM mechanism looks like it could provide an interesting way for new users to get a grid-ready UI by installing VirtualBox (or similar), downloading the CERN VM image, and applying a GridPP-specific context.

Site Updates

Tuesday 24th February

  • Next review of status today.

Tuesday 27th January

  • Squids not in GOCDB for: UCL; ECDF; Birmingham; Durham; RHUL; IC; Sussex; Lancaster
  • Squids in GOCDB for: EFDA-JET; Manchester; Liverpool; Cambridge; Sheffield; Bristol; Brunel; QMUL; T1; Oxford; Glasgow; RALPPD.

Tuesday 2nd December

  • Multicore status. Queues available (63%)
    • YES: RAL T1; Brunel; Imperial; QMUL; Lancaster; Liverpool; Manchester; Glasgow; Cambridge; Oxford; RALPP; Sussex (12)
    • NO: RHUL (testing); UCL; Sheffield (testing); Durham; ECDF (testing); Birmingham; Bristol (7)
  • According to our table for cloud/VMs (26%)
    • YES: RAL T1; Brunel; Imperial; Manchester; Oxford (5)
    • NO: QMUL; RHUL; UCL; Lancaster; Liverpool; Sheffield; Durham; ECDF; Glasgow; Birmingham; Bristol; Cambridge; RALPP; Sussex (14)
  • GridPP DIRAC jobs successful (58%)
    • YES: Bristol; Glasgow; Lancaster; Liverpool; Manchester; Oxford; Sheffield; Brunel; IC; QMUL; RHUL (11)
    • NO: Cambridge; Durham; RALPP; RAL T1 (4) + ECDF; Sussex; UCL; Birmingham (4)
  • IPv6 status
    • Allocation - 42%
    • YES: RAL T1; Brunel; IC; QMUL; Manchester; Sheffield; Cambridge; Oxford (8)
    • NO: RHUL; UCL; Lancaster; Liverpool; Durham; ECDF; Glasgow; Birmingham; Bristol; RALPP; Sussex
  • Dual stack nodes - 21%
    • YES: Brunel; IC; QMUL; Oxford (4)
    • NO: RHUL; UCL; Lancaster; Glasgow; Liverpool; Manchester; Sheffield; Durham; ECDF; Birmingham; Bristol; Cambridge; RALPP; Sussex, RAL T1 (15)

Tuesday 21st October

  • High loads seen in xroot by several sites: Liverpool and RALT1... and also Bristol (see Luke's TB-S email on 16/10 for questions about changes to help).

Tuesday 9th September

  • Intel announced the new generation of Xeon based on Haswell.

Meeting Summaries
Project Management Board - MembersMinutes Quarterly Reports


GridPP ops meeting - Agendas Actions Core Tasks


RAL Tier-1 Experiment Liaison Meeting (Wednesday 13:30) Agenda Meeting takes place on Vidyo.

Wednesday 11th March 2015

  • Operations report
  • Plan to swap out problematic Tier1 network router early tomorrow (Thursday 12th March) morning. (Delayed from last week).
  • Proposed date for Castor upgrade to version 2.1.14-15 is the 8th April. (To be confirmed).
  • The rollout of CVMFS version 2.1.20 is just about complete across the batch farm. As we are a test site for this we will check it runs OK for a further fortnight before reporting to CERN that it is OK.
  • All data has now been migrated off the T10000A, as well as the T10000B tapes.
WLCG Grid Deployment Board - Agendas MB agendas


NGI UK - Homepage CA


UK ATLAS - Shifter view News & Links

Atlas S&C week 2-6 Feb 2015


• Prodsys-2 in production since Dec 1st

• Deployment has not been transparent , many issued has been solved, the grid is filled again

• MC15 is expected to start soon, waiting for physics validations, evgen testing is underway and close to finalised.. Simulation expected to be broadly similar to MC14, no blockers expected.


• Rucio in production since Dec 1st and is ready for LHC RUN-2. Some fields need improvements, including transfer and deletion agents, documentation and monitoring.

Rucio dumps available.

Dark data cleaning

files declaration . Only Only DDM ops can issue lost files declaration for now, cloud support needs to fill a ticket.

• Webdav panda functional tests with Hammercloud are ongoing


Main page

DDM Accounting




• ASAP (ATLAS Site Availability Performance) in place. Every 3 months the T2s sites performing BELOW 80% are reported to the International Computing Board.





  • N/A
To note

  • N/A